Trouble with this question
If P(A) = O.41 then P(Ā)=
A) 0.59 B) 1.00 C) 0.26 D) 0.82

Answer by jorel1380(3719)

If P(A)=0.41, then P (Ā)=1 - P(A)=1-0.41=0.59
☺☺☺☺